5 years ago, bob was 2/3 mary's age. In 10 years, he will be 5/6 her age. How old is bob now?
a. 10 b. 15 c. 20 d. 21
How is this set up and how is it solved?

Answer by richwmiller(17219)

I'll give you the basics and let's see how you do.
Bob is age b now
Mary is age m now
5 years ago Bob was b-5 and Mary was m-5
In ten years Bob will be b+10 and Mary will be m+10
The biggest mistake that students make is forgetting to keep the ages balanced by adding or subtracting the same number form both.
I'll set up the first equation
5 years ago, bob was 2/3 mary's age
(b-5)=2/3*(m-5)
Now you set up the second equation.
In 10 years, he will be 5/6 her age
(b-5)=2/3*(m-5)
When you write the other equation, remember, they are asking for bob's age not mary's so solve for b
bob will be the younger because he is 2/3 of her age at one point and later 5/6 of her age.